The perimeter of a rectangle is 80 inches. The length of the rectangle is 8 less then three times the width. What are the dimensions of the rectangle?
Answer by checkley77(12844)

P=2L+2W
L=3W-8
80=2(3W-8)+2W
80=6W-16+2W
80=8W-16
8W=80+16
8W=96
W=96/8
W=12 ANS. FOR THE WIDTH.
L=3*12-8
L=36-8
L=28 ANS. FOR THE LENGTH.
PROOF:
80=2*28+2*12
80=56+24
80=80